# LASPATED: A Library for the Analysis of Spatio-Temporal Discrete Data (User Manual)

### v2

@@ 10 Replication script
Along with the LASPATED source code, we also provide a python script that performs all experiments from Guigues et al. (2023). Before running the script, the first step is to compile the C++ code. If you have a GCC compiler installed, this can be done by running the command:
@@
All results presented in the paper are saved in the directory replication_results, with subdirectories plots and tables. If you want access to the raw data generated by C++, check the directory cpp_tests/results, with subdirectories for each experiment made. Subdirectories ex1 and ex2 contain results for the first and second model without covariates (Examples 1 and 2). Subdirectory ex3 contains results for the model with covariates (Table 2 of the paper) and real_data contains results for the experiments with real data of Rio de Janeiro. For more details, see Guigues et al. (2023).
@@ 11 Docker container
We also provide a Docker container. The docker container comes with all dependencies installed and can be built in the project root directory using:
    \$ docker build -t laspated-dock .
and run with command:
    \$ docker run -it laspated-dock
If the user has a Gurobi WLS license, it can be added using:
    \$ docker run --volume="/absolute/path/to/gurobi.lic:/opt/gurobi/gurobi.lic:ro"
@@
This command will download all Python dependencies and install laspated on the \$PYTHONPATH, so you can use it in Python via:
    import laspated
See the demo directory for examples on how to run the discretization functions, and how to run the C++ calibration functions from Python.

### v3

@@ 10 Replication script
Along with the LASPATED source code, we also provide a python script that performs all experiments from laspatedpaper. Before running the script, the first step is to compile the C++ code. If you have a GCC compiler installed, this can be done by running the command:
@@
All results presented in the paper are saved in the directory replication_results, with subdirectories plots and tables. If you want access to the raw data generated by C++, check the directory cpp_tests/results, with subdirectories for each experiment made. Subdirectories ex1 and ex2 contain results for the first and second model without covariates (Examples 1 and 2). Subdirectory ex3 contains results for the model with covariates (Table 2 of the paper) and real_data contains results for the experiments with real data of Rio de Janeiro. For more details, see laspatedpaper.
@@ 11 Docker container
We also provide a Docker container. It is available at https://dockerhub.com or via the Dockerfile in the repository. The docker container comes with all dependencies installed and the library ready to use.
11.1 Dockerfile
To build the container via the repository Dockerfile, run:
    docker build -t laspated .
If you have a Gurobi Web License, you can build the container with Gurobi support by running:
    docker build --build-arg USE_GUROBI=1 -t laspated .
The above command will build the container with Gurobi 11.0.1 installed.
11.2 DockerHub
To download the container from DockerHub, just run:
    docker pull victorvhrn/laspated
Note that the DockerHub container is built without Gurobi support by default. To enable Gurobi support, the C++ code must be recompiled inside the container. Check the next section on how to run the container and section 8.2.6 on how to recompile the C++ code.
11.3 Running the container
To run the container, use:
    docker run -it laspated
To run the container with Gurobi support, you can pass a valid Gurobi Web license to the container with:
    docker run --volume="/absolute/path/to/gurobi.lic:/opt/gurobi/gurobi.lic:ro"
@@ 11.3 Running the container
Both commands will open a shell environment with all dependencies installed. Once in the container environment, you can run both the Python and C++ functions.
12 Missing Data
An example of model estimation with missing data is available in subdirectory Missing_Data. The subdirectory contains the implementation of four models that consider missing location data, and one artificial example considering missing timestamp data.
The file missing_data.hpp, implements the regularized model (class MissingLambdaRegularizedModel) and a model that assumes the probability of a missing location for an arrival is proportional to the population in each zone (class MissingLambdaCovariatesModel). Additionally, missing_data.hpp implements a customized version of the cross validation procedure for the regularized model.
The file missing_data.cpp implements a function that performs the estimation using a closed form expression, a function that uses the Fisher information matrix to compute confidence intervals, a function that runs an artificial example with missing timestamps, and a function that runs a set of experiments on the models. This code corresponds to the tests performed in the numerical experiments section of ref_missing_data.
Because of the Fisher information matrix computation, the missing data code requires the xtensor-blas and the openblas libraries. To compile the code, run:
    cd Missing_Data
    g++ -o missing Cpp/missing_data.cpp -DUSE_GUROBI=1 -std=c++14 -m64
    -I../Model_Calibration/Cpp -I\$GUROBI_HOME/include -L\$GUROBI_HOME/lib
    -lboost_program_options -lopenblas -O3 -lgurobi110 -lgurobi_c++
The signature of the cross validation function for missing data is given by:
laspated::CrossValidationResult cross_validation(
    laspated::Param& param, MissingLambdaRegularizedModel& model,
    xt::xarray<int>& sample, xt::xarray<int>& sample_missing_arrivals,
    std::vector<double>& group_weights)
The following example shows how to run the C++ code for missing data:
    ./missing -f test.cfg --model=all --info_file=info.dat
    --neighbors_file=neighbors.dat --arrivals_file=arrivals.dat
    --missing_file=missing_arrivals.dat
This example runs all missing data models for the input files provided. In general, options can be provided in a configuration file, the test.cfg file provided with -f, or in the command line. Note that the command line options override the values in the configuration file. The following options are identical to the LASPATED application:
  • EPS,
  • sigma,
  • max_iter,
  • lower_lambda,
  • beta_bar,
  • info_file,
  • arrivals_file,
  • neighbors_file.
The remaining options are specific to the missing data application. The option missing_file specifies the path to the file containing missing arrivals’ data. This file follows the same format as arrivals_file, but the application ignores the value of the zone index. Option mc_samples specifies the Monte Carlo sample size $S$, used in the MissingLambdaCovariatesModel. The option test_weights contains a list of weights to be used in the MissingLambdaRegularizedModel.
The option model defines the missing data model(s) that the application will run. It can be one of the following:
  • all: runs all four missing data models;
  • analytical: runs only the closed form models;
  • regularized: runs only the MissingLambdaRegularizedModel, once for each of the test_weights set, and once performing a cross validation;
  • population: runs only the MissingLambdaCovariatesModel, with the number of samples provided by mc_samples.
For the MissingLambdaCovariatesModel, there is also the function mean_var_u, that computes the mean and variance of the Monte Carlo samples for an observation indexed by $c\in\mathcal{C},d\in\mathcal{D},t\in\mathcal{T},n\in\mathcal{N}_{c,d,t}$. Its signature is:
    std::pair<double, double> mean_var_u(xt::xarray<int>& mn_samples, int c,
                                       int d, int t, int n,
                                       xt::xarray<double>& x)
where mn_samples is the $S\times|\mathcal{I}|$ array of samples, $(c,d,t,n)$ is the observation index and $x=(x_{c,d,t,i}:c\in\mathcal{C},d\in\mathcal{D},t\in\mathcal{T},i\in\mathcal{I})$ is the Poisson intensities array. The function returns a pair containing the mean and variance, respectively.